### v1

# MolBind: Multimodal Alignment of Language, Molecules, and Proteins

Abstract
Recent advancements in biology and chemistry have leveraged multi-modal learning, integrating molecules and their natural language descriptions to enhance drug discovery. However, current pre-training frameworks are limited to two modalities, and designing a unified network to process different modalities (e.g., natural language, 2D molecular graphs, 3D molecular conformations, and 3D proteins) remains challenging due to inherent gaps among them. In this work, we propose MolBind, a framework that trains encoders for multiple modalities through contrastive learning, mapping all modalities to a shared feature space for multi-modal semantic alignment. To facilitate effective pre-training of MolBind on multiple modalities, we also build and collect a high-quality dataset with four modalities, MolBind-M4, including graph-language, conformation-language, graph-conformation, and conformation-protein paired data. MolBind shows superior zero-shot learning performance across a wide range of tasks, demonstrating its strong capability of capturing the underlying semantics of multiple modalities. Our code and data can be found at MolBind.
1 Introduction
Figure 1: Illustration of modality data we try to align in this paper. The scientific language: text describes the name, property, and structure of molecules. 2D molecular graph: the atoms and bonds are nodes and edges in the graph. 3D molecular conformation: the atoms in the 3D Euclidean space. 3D protein pocket: a specific 3D region on the surface of a protein that can bind or interact with other molecules.
Multi-modal pre-training, as exemplified by the contrastive language-image pre-training model CLIP Radford et al. (2021), has made significant strides in computer vision and natural language processing. By aligning latent representations of images and textual descriptions, CLIP demonstrates remarkable zero-shot capabilities. This success spurred the adoption of multi-modal pre-training in many domains, including audio Guzhov et al. (2022), video Xu et al. (2021a), tabular data Hager et al. (2023), and point clouds Zhang et al. (2022).
Remarkably, the properties of molecules can be effectively described by language, i.e., biomedical texts with rich expressions. Consequently, multi-modal pre-training has been adopted in specialized scientific fields such as biology, facilitating advancements in molecule-language alignment Li et al. (2024); Su et al. (2022); Edwards et al. (2021). Molecule-language alignment has brought tremendous improvements for example for language-guided molecule generation Liu et al. (2023a); Zhu et al. (2024), molecule-language retrieval Edwards et al. (2021), and molecule-based text generation Liu et al. (2023b); Li et al. (2024).
Despite these successes, these approaches face two critical limitations: (i) The availability of multi-modal pairs for molecules is insufficient and significantly smaller compared to other domains. For example, CLIP is trained on a dataset of 400M image-language pairs collected from the internet, whereas the total number of available molecule-language pairs is only about 300K Liu et al. (2023a, b) which is orders of magnitude lower. Despite the achievements of multi-modal pre-training, its broader and more flexible application in biology domains remains limited by the high dependence on large-scale, high-quality paired data. (ii) Current multi-modal methods for molecules utilize only a single pair of modalities, limiting their extension to multiple N ($\geq 3$) modalities. Thus, the final representations are restricted to the pairs of modalities used during training. For instance, aligned molecule-language representations cannot be directly applied to conformation-protein tasks, and vice versa.
To address these challenges, we propose MolBind, a novel multi-modal pre-training framework that extends molecule-language pre-training to multiple modalities. Inspired by recent work Girdhar et al. (2023), attempting to bind various image-paired datasets, we bind all modalities (molecular language description, 2D molecular graph, 3D molecular conformation, 3D protein pocket) as shown in Figure 1 to a unified embedding space, enabling effective semantic alignment and facilitating a comprehensive understanding of various modalities in the biology domain. MolBind does not require datasets where all modalities co-occur for each sample, which are very hard to collect. Instead, it learns a single alignment representation space by leveraging multiple types of modality pairs. To validate our MolBind, we build and collect a high-quality dataset with 4 modalities, MoBind-M4 including graph-language, conformation-language, graph-conformation, and conformation-protein paired data. By augmenting the alignment process with data from multiple modalities, MolBind enhances model robustness against data sparsity and yields discernible improvements in downstream tasks.
Our key contributions are: (i) We propose a novel framework named MolBind, which effectively extracts shared semantic information from paired multi-modal data and gradually aligns various molecular modalities with the language modality into a joint space. (ii) We collect and construct a multi-modal dataset for molecules, comprising data pairs of four modalities. To the best of our knowledge, MoBind-M4 is the first unified dataset, containing multiple modalities of language, molecules, and proteins in the biological domain. (iii) Extensive experiments on downstream tasks, such as zero-shot cross-modal retrieval and cross-modal classification, validate the effectiveness of MolBind in representation alignment between language and molecular modalities.
2 Method
In this section, we present MolBind, a multi-modal pre-training method designed to align the latent semantic representations of multiple modalities in the biology domain with natural language.
2.1 Language, Molecule and Protein Encoders
In this section, we briefly introduce multi-modal encoders we choose. However. MolBind is conceptually simple and diverse models can be used as language, molecule, and protein encoders.
Language Encoder. The textual language modality delivers an extensive semantic analysis of the molecule’s structure, properties, and functionality. Such domain knowledge, in the form of natural language, is encoded using SciBERT Beltagy et al. (2019), which was pre-trained on textual data from the biological domain, following previous works Li et al. (2024); Liu et al. (2023b, a).
2D Graph Encoder. To encode the modality of 2D molecular graph, we utilize graph neural networks Hu et al. (2020); Xiao et al. (2022, 2021). Specifically, we use the pre-trained graph isomorphism network (GIN) encoder Hu et al. (2020) which is pre-trained on 2 million molecules from the ZINC15 Sterling and Irwin (2015) using graph contrastive learning You et al. (2020).
3D Conformation Encoder. We adopt Uni-Mol Zhou et al. (2022) as 3D conformation encoder (details in Appendix A.1). Specifically, Uni-Mol, a transformer-based model, has been pre-trained on 209 million molecular conformations and is capable of effectively capturing the input 3D geometry information of molecules.
3D Protein Encoder. Similar to 3D conformation, we also architect the 3D protein encoder as a Uni-Mol and initialize it from the checkpoint pre-trained on a 3M candidate protein pockets dataset.
We employ separate encoders for language, 2D molecular graphs, 3D molecular conformations, and 3D protein pockets. On each encoder, we add a modality-specific linear projection head to obtain a fixed-size, $D$-dimensional embedding. This embedding is normalized and utilized in subsequent contrastive learning to bind individual modalities.
2.2 Multi-modal Representation Alignment
MolBind utilize various pairs of modalities $(\mathcal{X},\mathcal{Y})$, where $\mathcal{X}$ represents one modality and $\mathcal{Y}$ another, to learn a single joint embedding. To ensure alignment across modalities, we adopt multi-modal contrastive learning principles Radford et al. (2021). Given a sample $x_{i}$ and its corresponding observation in the other modality $y_{i}$, we encode them into normalized embeddings: $\mathbf{x}_{i}=f_{\mathcal{X}}\left({x}_{i}\right)$ and $\mathbf{y}_{i}=f_{\mathcal{Y}}\left(y_{i}\right)$ where $f_{\mathcal{X}}$ and $f_{\mathcal{Y}}$ are encoders corresponding to modalities $\mathcal{X}$ and $\mathcal{Y}$. The encoders are optimized to increase the similarity of paired data, align them in the same semantic space, while minimizing the similarity of unpaired data:
$L_{\mathcal{X}2\mathcal{Y}}=-\log\frac{\exp(\mathbf{x}_{i}^{\top}\mathbf{y}_{i}/\tau)}{\exp(\mathbf{x}_{i}^{\top}\mathbf{y}_{i}/\tau)+\sum_{j\neq i}\exp(\mathbf{x}_{i}^{\top}\mathbf{y}_{j}/\tau)},$ (1)
where $\tau$ is the temperature and $y_{j}$ is the in-batch negative sample Radford et al. (2021). We use the symmetric loss to align $\mathcal{X}$ and $\mathcal{Y}$: $\mathcal{L}_{\mathcal{X}2\mathcal{Y}}+\mathcal{L}_{\mathcal{Y}2\mathcal{X}}$. MolBind applies this loss framework to jointly bind 4 pairs of modalities, including molecular language-graph, language-conformation, graph-conformation, and conformation-protein.
Table 1: Performance of cross-modal retrieval on test sets of MolBind-M4 includes the following scenarios: Graph to Language (G2L), Language to Graph (L2G), Language to Conformation (L2C), and Conformation to Language (C2L) retrieval.
Retrieval in batch Retrieval in full test set
G2L (%) L2G (%) C2L (%) L2C (%) G2L (%) L2G (%) C2L (%) L2C (%)
Model R@1 R@20 R@1 R@20 R@1 R@20 R@1 R@20 R@1 R@20 R@1 R@20 R@1 R@20 R@1 R@20
Sci-BERT 78.2 89.5 76.1 88.5 77.2 88.8 77.9 88.3 61.0 85.2 55.7 86.3 52.6 84.7 49.0 91.2
KV-PLM 78.5 90.3 76.6 89.1 77.9 89.2 78.8 88.7 61.5 85.9 56.1 86.6 53.2 85.1 49.3 91.5
MoMu 80.1 91.5 78.4 90.2 79.1 90.5 79.3 91.1 63.9 87.6 58.0 87.8 54.8 86.3 50.2 93.1
MolCA 85.3 94.2 82.5 94.8 83.9 94.7 83.5 94.2 67.1 91.8 60.5 91.2 58.2 90.7 55.6 95.6
MolBind 88.5 98.6 86.5 98.7 86.0 98.6 85.1 98.5 71.5 93.0 64.3 93.6 62.4 94.0 58.9 98.5
2.3 Dataset Collection
In this section, we briefly describe how to construct the MolBind-M4 dataset (full details in Appendix A.2), including 325K pairs of language-graph data, 161K pairs of language-conformation data, 161K pairs of graph-conformation data, and 72K pairs of conformation-protein data.
Language-Graph Pairs: PubChem Kim et al. (2023), the most extensive source of chemical information, provided many molecules with diverse descriptions. We transformed these molecules into graphs and paired them with corresponding textual descriptions, yielding 32K language-graph pairs.
Language-Conformation Pairs: Using Molecule3D Xu et al. (2023), which contains 3.9M ground-state molecular conformations, and the GEOM dataset Axelrod et al. (2022) with 37M higher-quality conformations, we matched molecular IDs (CIDs) and InChIs in them with textual descriptions from PubChem. This process resulted in 157K language-conformation pairs.
Conformation-Protein Pairs: By leveraging data from PDBBind Wang et al. (2005) and CrossDocked Francoeur et al. (2020), we extracted and refined conformation-protein pairs. After filtering based on RMSD thresholds and clustering for sequence identity, we consolidated our findings into 72K unique pairs, ensuring no overlap with previously identified pocket proteins.
Graph-Conformation Pairs: Converting 3D molecular structures from the language-conformation dataset into 2D graphs allowed us to generate 158K graph-conformation pairs.
3 Experiments
In this section, we evaluate the effectiveness of MolBind in various downstream tasks through different experiments. Firstly, MolBind’s capability to align molecular modalities and text is evaluated using zero-shot cross-modal retrieval and classification. In addition, we use MolBind to enhance the performance of downstream tasks that involve molecules and proteins. The pre-training settings can found in Appendix A.3.
3.1 Zero-shot Molecule-Language Retrieval
We evaluate the graph-language and conformation-language retrieval performance on the MoBind-M4 dataset. For our baselines, we employ Sci-BERT Beltagy et al. (2019), KV-PLM Zeng et al. (2022), MoMu Su et al. (2022), and MolCA Liu et al. (2023b), assessing their performance using Recall@1 and Recall@20. These metrics are evaluated both within a batch of 64 samples and across the entire test set, following the methodology outlined in Liu et al. (2023b). As presented in Table 1, MolBind’s performance surpasses that of MolCA by a large margin. Overall, MolBind significantly outperforms prior work on two molecule-language retrieval tasks, validating its effectiveness in aligning the molecule and language modalities.
Figure 2: Zero-shot molecule classification with language.
3.2 Zero-shot Classification with Language
The International Union for Pure and Applied Chemistry (IUPAC) has established a standardized naming system for chemical compounds, known as IUPAC names Favre and Powell (2013). Notably, this system relies on identifying specific molecule structures. Therefore, correctly predicting IUPAC names indicates a model’s ability to understand molecule structures. MolBind has the capability for zero-shot IUPAC namultiple modalityn, thanks to its aligned representation space of molecule and language descriptions (see detailed implementation in Appendix A.4). For evaluation, we collect IUPAC names for molecules in valid and test sets of MolBind-M4 using the PubChemPy library, and report average accuracy of test set. We compare the following baselines: MoMu Su et al. (2022), MolCA Liu et al. (2023b), Text2Mol Edwards et al. (2021) and MoleculeSTM Liu et al. (2023a).
Figure 2 shows our MolBind consistently outperforms baselines by a large margin, showing that modeling multiple modality for molecules is an important factor in comprehending molecule structures. These results also demonstrate the data efficiency of MolBind-induced zero-shot classifiers based on the aligned representation space of 2D graph and language descriptions.
3.3 Zero-shot Molecule to Protein Retrieval
Since MolBind has the ability to align the representations of molecular conformations and proteins, it can also be used for conformation-to-protein retrieval, which is known as target fishing in drug discovery. We established a benchmark using the CASF-2016 dataset. We compared MolBind with Vina Varadi et al. (2022), Glide Halgren et al. (2004), and DrugCLIP Gao et al. (2023). As shown in Table 2, our MolBind significantly outperforms Vina, Glide, and DrugCLIP. More importantly, the improvement of MolBind over DrugCLIP, which utilizes only two modalities, validates its ability to align molecular conformations and protein modalities using multiple modalities.
3.4 Ablation Study
In Table 3, we analyze the effect of different modalities in MolBind’s multi-modal pretraining. When each objective function is individually applied (1st, 2nd and 3rd row), the modality for (language, graph) performs best on L2G, which probably benefits from its explicit alignment. Interestingly, we observe an emergent behavior in the embedding space that aligns two pairs of modalities (language, conformation) even though we only train using the pairs (language, graph) and (graph,conformation). These results demonstrate that MolBind aligns the modalities and implicitly transfers the language supervision associated with conformation to other modalities such as graph. Overall, MolBind achieves the best performance, validating that it is an effective way to align multiple modalities simultaneously.
Table 2: Conformation-to-protein retrieval on CASF-2016.
Conformation to Protein (Recall)
Model R@1 R@2 R@3 R@4 R@5
Vina 3.38 5.26 7.52 9.02 10.15
Glide 14.98 22.85 30.34 35.58 39.33
DrugCLIP 24.07 42.96 51.11 59.26 62.59
MolBind 45.96 48.31 60.15 68.25 75.92
4 Conclusions and Discussion
Table 3: Ablation study for using different modalities on language to graph (L2G) and language to conformation (L2C) retrieval tasks. We follow the same setting in Section 3.1.
Ablations L2G L2C
(language, graph) 60.4 0.23
(language, conformation) 0.51 55.7
(graph, conformation) 0.25 0.31
(language, graph) & (language, conformation) 62.8 58.1
(language, graph) & (graph, conformation) 61.4 56.1
(language, conformation) & (graph, conformation) 60.4 57.6
MolBind 64.3 58.9
MolBind is a simple yet effective framework that aligns multiple modalities related to molecules in a joint embedding space. MolBind achieves multi-modal alignment under the constraints of insufficient biochemical data, facilitating knowledge transfer between modalities. We also release the multi-modal dataset MolBind-M4, which is the first dataset to combine multiple modalities, sourced exclusively from openly accessible datasets. We evaluate MolBind in various downstream tasks, demonstrating its effectiveness in aligning representations from multiple modalities.
5 Limitations and Impact Statements
This paper presents work aimed at learning a unified joint embedding space for multiple modalities related to language and molecules. While we provide a high-quality, multi-modal training dataset, the scale of our data is comparatively smaller than that in the vision-language domain. Expanding and refining multi-modal pre-training datasets is a crucial future work for scientific research scenarios.
Appendix A Appendix
A.1 The Details of Uni-Mol Encoder
In this section, we provide additional details about the 3D encoder in Uni-Mol Zhou et al. (2022). We first learn the atom embedding which is initialized using an embedding matrix based on the input atomic features of molecules $\mathbf{h}$:
$\mathbf{m}=\mathbf{h}\mathbf{W}_{h},$ (2)
where $\mathbf{W}_{h}\in\mathbb{R}^{d\times D}$ is the learnable parameter matrix and $D$ is the hidden dimension. With the atom embedding, Uni-mol initially learns the pair-type-aware Gaussian kernel Shuaibi et al. (2021) to model pairwise representation for atom pairs. This process is denoted as follows.
$\mathbf{p}_{ij}={\mathcal{G}(\mathcal{A}\left(d_{ij},t_{ij};\mathbf{a},\mathbf{b}\right),\mu^{k},\sigma^{k})\mid k\in[1,D]},$ (3)
$\mathcal{A}(d,r;\mathbf{a},\mathbf{b})=a_{r}d+b_{r},$
where $d_{ij}$ is the Euclidean distance between coordinates $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ for the atom pair $ij$, $\mathcal{G}(d,\mu,\sigma)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(d-\mu)^{2}}{2\sigma^{2}}}$ is a Gaussian density function with parameters $\mu$ and $\sigma$. $\mathcal{A}\left(d_{ij},t_{ij};\mathbf{a},\mathbf{b}\right)$ is the affine transformation with parameters $\mathbf{a}$ and $\mathbf{b}$. This function affines $d_{ij}$ corresponding to its pair-type $t_{ij}$.
The pair representation is utilized to enhance the 3D spatial encoding within the model Zhou et al. (2022). It is updated via the multi-head Query-Key product results in self-attention in the layer $l$:
$\mathbf{q}_{ij}^{l+1}=\mathbf{q}_{ij}^{l}+\{\frac{\mathbf{h}_{i}^{l}\mathbf{W}_{l,h}^{Q}(\mathbf{h}_{j}^{l}\mathbf{W}_{l,h}^{K})^{T}}{\sqrt{D}}\rvert\,h\in[1,H]\},$ (4)
where $\mathbf{h}_{i}^{l}$ is the atom representation of the $i$-th atom at $l$-th layer and $\mathbf{q}_{ij}^{l}$ is the pair representation matrix of atom pair $ij$ in $l$-th layer. $H$ is the number of attention heads, $D$ is the dimension of hidden representations. $\mathbf{W}_{l,h}^{Q}$ and $\mathbf{W}_{l,h}^{K}$ are the learnable parameter matrix of the projection for Query (Key) of the $l$-th layer $h$-th head. $\mathbf{q}_{ij}^{0}$ is the input of the model, which is initialized with $\mathbf{p}_{ij}$.
Following the acquisition of the pair representation, it is used by a self-attention mechanism to get the atom representations:
$\mathbf{h}_{i}^{l+1,h}$ $=\operatorname{softmax}(\frac{\mathbf{h}_{i}^{l}\mathbf{W}_{l,h}^{Q}(\mathbf{h}_{j}^{l}\mathbf{W}_{l,h}^{K})^{T}}{\sqrt{d}}+\mathbf{q}_{ij}^{l,h})\mathbf{h}_{j}^{l}\mathbf{W}_{l,h}^{V},$ (5)
$\mathbf{h}_{i}^{l+1}$ $=\operatorname{concat}_{h}(\mathbf{h}_{i}^{l+1,h}),$
where $\operatorname{softmax}(*)$ is the softmax function, $\operatorname{concat}_{h}(*)$ is the concatenation of the representation from different heads and $\mathbf{h}_{i}^{l+1,h}$ is the representation vector for the atom $i$ of head $h$ in the layer $l+1$. $\mathbf{W}_{l,h}^{V}$ is the learnable parameter matrix, which is the projection of Value of the layer $l$ and head $h$. The representation of the first layer $\mathbf{h}_{i}^{0}$ is initialized with the atom embedding $\mathbf{m}$. We denote the final output as $\mathbf{h}_{i}^{L}$ (where $L$ represents the number of layers) as $\mathbf{z}_{i}$, and the final representation from the 3D encoder is obtained after the mean pooling of all atom representations.
Table 4: Statistics and details of MolBind-M4. We randomly sample subsets for validation and testing across all modalities except for (conformation, protein), which has official validation and test sets (CASF-2016).
Modality Pretrain Validation Test Data Sources
(language, graph) 319,353 1,500 1,500 PubChem
(language, conformation) 158,237 1,500 1,500 Molecule3D, GEOM
(conformation, protein) 72,355 100 285 PDBBind, CrossDock
(graph, conformation) 158,237 1,500 1,500 Molecule3D
A.2 Details of Constructing MolBind-M4
There are various modalities in biology and chemistry such as language, 2D graph, 3D conformation, 3D protein (large biomolecules). In light of the varied origins of the data and potential high computational and expert annotation costs, attaining a large amount of comprehensive molecular data across all modalities is impractical. Previous studies Girdhar et al. (2023); Zhu et al. (2023) have shown that modalities can be aligned via intermediary modality even without explicitly paired data.
Given the presence of numerous challenging-to-obtain data, especially in the field of biochemistry, such as annotations and high-quality molecular conformations, paired alignment methods are heightened. More notably, using this approach can also conveniently facilitate knowledge transfer among modalities, contributing to learning representations for modality that suffer from sparsely available data. To this end, we collected and curated four types of cross-modal molecule related pairs all from open sources to learn a single joint embedding space. In the following, we detail the construction of our MolBind-M4, the statistics and data source of which are provided in Table 4.
Language-Graph Pairs. We used Pug View Kim et al. (2019) to download molecule annotations. In total, we obtained 349,356 molecules with textual descriptions. Most of these annotations begin with the common name or the International Union of Pure and Applied Chemistry (IUPAC) name Kuhn et al. (2004). To encourage the text encoder to learn generalized descriptions of molecule functionality and properties, we uniformly replaced descriptions that start with a name with ’The molecule.’ We sampled high-quality validation and test sets, each consisting of 1,500 pairs with texts longer than 19 words. The remaining samples were allocated to a pre-training set, totaling 319,353 samples.
Language-Conformation Pairs. Although tools like RDKit Bento et al. (2020) enable the generation of molecular conformations using methods such as ETKDG Riniker and Landrum (2015) and empirical force fields Rappé et al. (1992), these approaches often suffer from inaccuracies and poor generalizability Volokhova et al. (2023). Thus, we consider ground-state conformations.
Specifically, we initially acquired molecular conformations from two sources. The first dataset, Molecule3D Xu et al. (2021b), includes 39M ground-state molecular conformations computed via DFT from PubChemQC Nakata and Shimazaki (2017). These conformations were then matched with corresponding textual descriptions in language-graph pairs using CID as a unique identifier, resulting in 25, 961 conformation-language pairs. The second dataset, GEOM, contains approximately 37M conformations generated using a semi-empirical DFT algorithm. Due to the absence of CIDs, we matched conformation and language-graph pairs using the International Chemical Identifier (InChI) Heller et al. (2015) to obtain textual descriptions and identify molecules. Through this process, we obtained a total of 158, 327 high-quality conformations. For data partitioning, we sampled 1,500 samples for both the training and testing sets from the most accurate ground-state conformations derived from Molecule3D. During the sampling process, we ensured that the text did not appear in the training set of language-graph pairs to mitigate potential data leakage.
Conformation-Protein Pairs. We initially extracted 19, 443 experimentally validated molecule-protein structure pairs from PDBBind Wang et al. (2005). We then expanded our dataset with the CrossDocked dataset, where structures were generated by docking ligands and pockets from the same class of binding site as delineated by Pocketome Kufareva et al. (2012), using smina. Following previous work Luo et al. (2021); Peng et al. (2022), we filtered out all poses with RMSD greater than 1Å and used MMseqs2 to cluster the data at 30% sequence identity. We randomly selected 100K protein-ligand pairs for pre-training and another 100 pairs for the validation set. Afterward, we merged these 100K pairs with PDBBind, excluding any duplicates to form the final pre-training set, which contains 72,355 pairs. Finally, we used the widely recognized CASF2016 dataset from PDBBind as the testing dataset, which includes 285 experimentally validated structures.
Graph-Conformation Pairs. Given that 3D molecules can be converted to 2D representations using chemical tools like RDKit, we directly processed molecular structures from Conformation-Language pairs, converting them into 2D graphs. We preserved the same dataset split, incorporating 158,237 graph-conformation pairs into the pre-training set and splitting 1,500 pairs each to the training and testing sets.
A.3 Multi-modal Pre-training Details
We train our model using the Adam optimizer with a learning rate of 0.001. The other hyper-parameters for encoders are set to their default values. We have a batch size of 16, and we use 4 NVIDIA A100 GPU cards for acceleration. We train our model for a maximum of 100 epochs. The temperature is set to 1 and we find that it has less effect on the performance. For more detailed training configurations, please refer to our code.
A.4 Details for Zero-shot Molecule Name Classification with Language
Given the 2D graph $G$ of the query molecule and language “$\text{\textquotedblleft The molecule is }\{\text{IUPAC name}\}."“$ with the IUPAC name $\left\{t_{i}\right\}_{i=1}^{M}$ of all $M$ names, we employ the graph encoder to extract molecule representation $\mathbf{g}$ and use the jointly learned text encoder to extract textual name representations $\left\{\textbf{t}_{i}\right\}_{i=1}^{M}$. We then derive classification logits $\left\{y_{i}\right\}_{i=1}^{M}$ by comparing the dot product similarity between molecule and name representations: $y_{i}=\mathbf{g}\cdot\textbf{t}_{i}/\tau(i=1,\cdots,M)$, which follows the formula of the InfoNCE loss in Equation 1 in the main paper. Softmax is performed on these logits to derive classification probabilities over IUPAC name space.

### v2

@@
 2.3 Dataset Collection
-In this section, we briefly describe how to construct the MolBind-M4 dataset (full details in Appendix A.2), including 325K pairs of language-graph data, 161K pairs of language-conformation data, 161K pairs of graph-conformation data, and 72K pairs of conformation-protein data.
+In this section, we briefly describe how to construct the MolBind-M4 dataset (full details in Appendix A.2), including 322K pairs of language-graph data, 161K pairs of language-conformation data, 161K pairs of graph-conformation data, and 72K pairs of conformation-protein data.
 Language-Graph Pairs: PubChem Kim et al.
 (2023), the most extensive source of chemical information, provided many molecules with diverse descriptions.
-We transformed these molecules into graphs and paired them with corresponding textual descriptions, yielding 32K language-graph pairs.
+We transformed these molecules into graphs and paired them with corresponding textual descriptions, yielding 322K language-graph pairs.
 Language-Conformation Pairs: Using Molecule3D Xu et al.
@@
 (2022) with 37M higher-quality conformations, we matched molecular IDs (CIDs) and InChIs in them with textual descriptions from PubChem.
-This process resulted in 157K language-conformation pairs.
+This process resulted in 161K language-conformation pairs.
 Conformation-Protein Pairs: By leveraging data from PDBBind Wang et al.
@@
 After filtering based on RMSD thresholds and clustering for sequence identity, we consolidated our findings into 72K unique pairs, ensuring no overlap with previously identified pocket proteins.
-Graph-Conformation Pairs: Converting 3D molecular structures from the language-conformation dataset into 2D graphs allowed us to generate 158K graph-conformation pairs.
+Graph-Conformation Pairs: Converting 3D molecular structures from the language-conformation dataset into 2D graphs allowed us to generate 161K graph-conformation pairs.
 3 Experiments
@@
 (2019) to download molecule annotations.
-In total, we obtained 349,356 molecules with textual descriptions.
+In total, we obtained 322,353 molecules with textual descriptions.
 Most of these annotations begin with the common name or the International Union of Pure and Applied Chemistry (IUPAC) name Kuhn et al.
@@
 (2015) to obtain textual descriptions and identify molecules.
-Through this process, we obtained a total of 158, 327 high-quality conformations.
+Through this process, we obtained a total of 158, 237 high-quality conformations.
 For data partitioning, we sampled 1,500 samples for both the training and testing sets from the most accurate ground-state conformations derived from Molecule3D.